A mass breaks suddenly into two parts, masses M and m, with speeds V and v respectively and carrying a total kinetic energy K. If (M/m) = R = 4.2 what fractions of KK does each of the masses carry? (Think about significant figures.)
1
Expert's answer
2020-03-25T11:05:40-0400

From the conservation of momentum:


"mv=MV=RmV\u2192v=RV"

"K=0.5mv ^
2
 +0.5Rm\frac{v^2}{R^2}=0.5mv ^2\left(1+\frac{1}{R}\right)"

"K_m=\frac{1}{(1+\frac{1}{R})}K=\frac{1}{(1+\frac{1}{4.2})}K=0.81K"

"K_M=K-K_m=(1-0.81)K=0.19K"